OpenStudy (anonymous):

derivative of e^(9x)=sin(x+7y)

OpenStudy (anonymous):

\[9e^{9x}=\cos(x+7y)\times \left(1+7y'\right)\] solve for \(y'\)

OpenStudy (anonymous):

can you help me through this part im not sure

OpenStudy (anonymous):

the hard part is knowing how to use the chain rule properly to get started with what i wrote above is that part ok?

OpenStudy (anonymous):

the chain rule is first times derivative of the second plus second times derivative of the first right?

OpenStudy (anonymous):

no, that is the product rule

OpenStudy (anonymous):

oh ok which is the chain rule again?

OpenStudy (anonymous):

\[\left(f(g)\right)'=f'(g)g'\]

OpenStudy (anonymous):

for example \[\frac{d}{dx}e^{9x}=e^{9x}\times \frac{d}{dx}9x=9e^{9x}\]

OpenStudy (anonymous):

and \[\frac{d}{dx}\sin(x+7y)=\cos(x+7y)\times \frac{d}{dx}[x+7y]\] \[=\cos(x+7y)\times (1+7y')\]

OpenStudy (anonymous):

9e^(9x)(9x^(-1/2))=cos(7x+14y)